# Weather-Driven Agricultural Decision-Making Using Digital Twins Under Imperfect Conditions

### v1

@@ 1. Introduction
Modern agriculture, especially high-value specialty crop production, relies heavily on data-driven automation technologies for production management. Agricultural decision-making increasingly relies on precise and consistent environmental data to optimize crop health, resource utilization, and yield predictions. Regional as well as localized weather critically drives several on-farm decision-support models and intelligent automated systems, including irrigation, abiotic stress (e.g., cold, heat) mitigation, and pest management, among others (boomgard2022machine, ). Weather data measurements, which include essential metrics such as temperature, humidity, and precipitation, form the foundation of such decisions. However, weather networks, like any cyber-physical and Internet-of-Things (IoT) systems, are often plagued by inconsistencies [^1] In this work, we focus on effects on inconsistent/irregular weather streams rather than the causes that trigger the anomalies. due to sensor faults, calibration drift, communication errors, environmental interference, or cyber breaches. Imperfect measurements can significantly undermine the reliability of downstream applications in agriculture, ranging from predictive modeling to real-time interventions. Such discrepancies can lead to inaccurate predictions, suboptimal resource allocation, and reduced agricultural productivity.
Any inconsistent weather data input could lead to inadequate decision-making and wrongful actuation of precision management scenarios, resulting in considerable crop yield and quality losses. To address this problem, we adopt the digital twin technology (barricelli2019survey, ) for analyzing and understanding agricultural weather networks that are critical for grower decision support. We build dynamic virtual data streams that mirror physical stations—to monitor, validate, and enhance the reliability of weather data in the face of anomalies. While digital twin technology is used in other cyber-physical domains (e.g., control systems, manufacturing, robotics), its benefits are not fully explored for digital agriculture. By leveraging the concept of digital twins, we show that it is feasible to create a virtual replica of weather networks for real-time inconsistency detection, data imputation, and predictive analytics. We name our framework Cerealia. [^2] In ancient Roman religion, the Cerealia was a major festival dedicated to the sowing, growing, and harvesting of crops. As our work aims to assist in better decision-making for agricultural stakeholders, we name our framework Cerealia.
@@
Cerealia is implemented and tested by using a commercial orchard located in Central Washington, United States (§4–§5). [^3] This research uses weather stations installed in a commercial orchard maintained by Washington State University, United States. While this is a closed system, access can be made available for research purposes by request, as we did in this work. We use this platform to demonstrate the efficacy of our approach in a real, live setup. Our use of this platform, however, by no means implies our affiliation with this infrastructure/institution. The digital twin modules are deployed on an NVIDIA Jetson Orin platform (karumbunathan2022nvidia, ). In addition to live weather feed from the orchard, Cerealia is also tested using a publicly available European weather dataset (jena_weather_dataset, ). We conduct a comprehensive evaluation of nine state-of-the-art neural network models to analyze their effectiveness in detecting weather data inconsistencies. We measure the overheads of running learning models on our evaluation platform and test the scalability of Cerealia. We find that the overheads of checking (in terms of timing and memory usage) are minimal (less than 1 s and 0.5 MB of memory use). Our research highlights the potential of digital twin tools like Cerealia to improve agricultural decision-making under imperfect conditions.
Cerealia implementation is publicly available: https://anonymous.4open.science/r/ag-dt-C7BC.
@@ 2.1. Weather Data Impacts on Agriculture
Weather data has a significant impact on agriculture. For example, in early spring and late winter, frost damage in fruit crops can hurt fruit production. Fruit growers attempt to mitigate frost using wind machines, propane heaters, and under-tree sprinklers (evans1999frost, ). To activate those mitigation tools, growers use weather data-driven models and sensing systems that automate and improve frost management with resource conservation (natural gas/energy) (talsma2023frost, ; cann2022observing, ). Similar to frost, extreme and more frequent heat stress events in summer can cause fruit cullage due to heat stress instigated by sunburn and related damages. To address this problem, researchers have explored weather-driven models to increase the effectiveness of heat stress mitigation techniques (goosman2023apple, ; ranjan2020field, ; amogi2023mask, ). In addition to frost and heat management, uninterrupted and reliable weather data obtained at the farm scale is necessary for several real-time tasks, including irrigation, disease and pest management, fruit growth prediction, among others (boomgard2022machine, ).
Existing weather-driven agriculture decision-making models (talsma2023frost, ; cann2022observing, ; goosman2023apple, ; ranjan2020field, ; saxena2023grape, ; amogi2023mask, ; amogi2024edge, ) rely on the availability of “perfect” weather information. Like any networked cyber-physical system, weather stations are also susceptible to faults and cyberattacks (boomgard2022machine, ; weiwei2018fault, ; botschner2024cybersecurity, ; kulkarni2024review, ). Inadequate and inconsistent environmental data, for instance, due to faults in weather station sensors or adversarial breaches, often fail to protect the crop during extreme heat or cold events. Further, any unreliable decision-making model or lack of automation requires the growers to work on the farm in extreme weather conditions to manually activate protection measures (e.g., turning on heathers, putting protective netting, or tuning on the cooling system to prevent heat stress). We envision that a digital twin platform incorporating measurement inconsistencies provides a virtual view of the weather network and can help growers make informed decisions in their target agriculture applications.
@@ 2.2. Digital Twin and its Usefulness in Agriculture
In literature, a digital twin is defined as “a digital replica of a living or non-living physical entity” (el2018digital, ). Digital twin technology is used to describe and model the corresponding physical counterpart digitally by combining modeling, simulation, and emulation technologies with other analytics to better understand aspects of the target application domain. Digital twins are commonly used in many cyber-physical domains such as manufacturing (kritzinger2018digital, ), IoT (minerva2020digital, ), healthcare (sun2023digital, ), and recently, digital agriculture (purcell2023digital, ). However, its efficacy in weather-driven agricultural decision-making still has not been explored.
@@ 3.1.1. Physical Layer
As shown in Block 1 of Fig. 2, the physical layer consists of weather stations deployed in the field. Many off-the-shelf weather stations (e.g., Cabled Vantage Pro2 (davisinstrumentsCabledVantage, ), ATMOS 41W (metergroupATMOSMETER, )) and existing weather networks (e.g., AgWeatherNet (awn, )) generally collect field measurements from stations in a periodic interval and store them in a cloud gateway for further processing. We follow a similar approach in Cerealia, where a cloud storage service collects data from distributed stations and feeds this to our analysis engines in the digital twin layer (see next).
@@ 3.1.2. Digital Twin Layer
The key components of the digital twin layer (i.e., Blocks  3- 7) are modular by design. As mentioned above, our current setup incorporates four kinds of noisy data and uses nine machine-learning models to see the models’ behavior under these inconsistencies to find the best possible solution for Cerealia. However, other faulty/anomalous data and other statistical models or anomaly detection tools can be incorporated with Cerealia to check for inconsistencies. Likewise, as discussed in §6, for demonstration purposes, we tested Cerealia for two agricultural use cases (e.g., weather data imputation and fruit surface temperature prediction problems). However, this component of Cerealia (i.e., Block  7) can be adapted for any other agriculture decision-making tasks (e.g., fruit bloom phenology (chaves2015modeling, ), soil water content prediction (nath2024soil, )) without loss of generality.
@@ 3.2. Generating Noisy Measurements
One of the benefits of having a digital twin architecture like Cerealia is that we can emulate the behaviors of actual weather stations with customized parameters. Although physical station measurements could be anomalous in practice, (a) they are unpredictable, and (b) we need a reference model (i.e., ground truth) to check for abnormalities. Hence, we use a noise generator (de2016benchmark, ) to mimic inconsistent behaviors. We consider four kinds of inconsistent data: (a) random (random spikes in the measurements), (b) malfunction (abrupt oscillations in the measurements), (c) drift (similar to malfunction, but the results are significantly shifted from nominal values), and (d) bias (results are scaled up or down than normal values). Figure 3 illustrates these four kinds of faults. Although we choose these four kinds of inconsistencies for our current demonstration—as they often represent typical sensor faults (ni2009sensor, )—any other fault model can be integrated with Cerealia’s synthetic data generator module (i.e., Block 3 in Fig. 2), which then will be fed into the inconsistency checker (i.e., Block 5 in Fig. 2, discussed next).
@@ 3.3. Leveraging Machine Learning for Inconsistency Checking
Based on historical traces and a given noise model, we train machine learning models to characterize off-nominal data from real-time weather measurements generated from the physical layer. Some machine learning models, such as SVM (suthaharan2016support, ), KNN (steinbach2009knn, ), and K-Means (steinley2006k, ), often assume static data relationships or require handcrafted features, which may not be efficient for capturing the complex sequential nature of weather measurements. Also, they often struggle with large datasets, particularly high-dimensional and noisy data (which is the case for weather networks). Hence, we use deep learning models in Cerealia to check for inconsistencies. In our current implementation, we tested nine neural network models, which are discussed next. However, as mentioned earlier, other statistical or learning models can be integrated with Cerealia.
@@ 3.3.1. Convolutional Neural Network (CNN) Models
CNN models stand out in retrieving hierarchical and local features from time-series data like weather measurements. We use two CNN models: (a) Temporal Convolution Network (TCN) (he2019temporal, ) and (b) ResNet (he2020resnet, ). TCN uses causal convolution and dilation to capture short-term and long-term temporal dependencies. ResNet uses residual connections and expedites the training of deep networks with reduced training error, but often requires careful tuning due to its complexity.
@@ 3.3.2. Recurrent Neural Network (RNN) Models
RNNs are highly effective for sequential data analysis due to their inherent memory capabilities. RNN models can capture the features and dependencies across long sequences. We incorporate the following RNN models for inconsistency checking: (a) Long Short-Term Memory (LSTM) (van2020review, ), (b) Bi-LSTM (siami2019performance, ) and (c) Gated Recurrent Unit (GRU) (shewalkar2019performance, ). LSTM is a well-known algorithm for retaining information over time. The latter two models are improved variants of LSTM that enhance sequence prediction.
@@ 3.3.3. Transformer Models
Transformer architectures are useful in tracking relationships in sequential data. We use two transformer-based models: (a) Time-Series Transformer (TST) (zerveas2021transformer, ) and (b) Informer (zhou2021informer, ) that are useful for capturing intricate patterns and long sequences in the weather station measurements.
@@ 3.3.4. Hybrid Models
We also use two hybrid models: (a) TST-AE (zhang2021unsupervised, ) and (b) LSTM-AE (provotar2019unsupervised, ). TST-AE uses a transformer-based self-attention with an encoder-decoder structure for better feature extraction. LSTM-AE is based on an LSTM-based encoder-decoder and can provide efficient memory usage, but comes with the cost of reduced accuracy.
@@ 4.1. Physical Layer Integration and Real-Time Data Collection
We integrate Cerealia with a commercial apple orchard located in Quincy, Washington, United States ($47^{\circ}13^{\prime}31.3^{\prime\prime}$ $119^{\circ}57^{\prime}38.0^{\prime\prime}\text{W}$). The orchard has an average elevation of 322 m and an average slope of $2^{\circ}$ facing east. The orchard’s weather stations are maintained by the Washington State University AgWeatherNet (awn, ) team. [^5] Note: The Majority of commercial agricultural weather networks are proprietary. Access to orchards/weather data is non-trivial, and few are publicly accessible for research. The station where we have access is installed at 1.5 m above ground level. The station uses an all-in-one weather sensor (ATMOS 41 from METER Group (metergroupATMOSMETER, )). It measures several weather parameters such as solar radiation, precipitation, vapor pressure, relative humidity, barometric pressure, horizontal wind speed, wind gust, wind direction, tilt, lightning strike, and lightning average distance.
Weather stations sample measurements at 5-minute intervals and send the data to a vendor cloud service (ZENTRA Cloud (metergroupZENTRACloud, )). We obtained weather data from March 2023 to January 2025. The collection frequency is twelve data tuples per hour per station (recall: stations send data to the cloud in 5-minute intervals), resulting in 192,613 data points. Following the naming convention suggested by the AgWeatherNet team, we refer to the measurements obtained from the stations as the “Quincy” network data. Our physical layer implementation demonstrates the feasibility of integrating Cerealia with a live operational weather network.
@@ 4.1.1. Preexisting Dataset
In addition to the live feed obtained from the orchard (Quincy network), we also use a publicly available dataset (named “Beutenberg”) (jena_weather_dataset, ). This weather data was recorded at the weather stations of the Max Planck Institute for Biogeochemistry in Jena, Germany. The dataset contains 20 different features, such as air temperature, atmospheric pressure, and humidity, between December 2015 to December 2020. The sampling rate for this dataset is six data points per hour. The total number of data points is 266,609. The reason for selecting this dataset is to test our inconsistency classifiers for diverse weather conditions (North America and Europe). Table 7 in Appendix A.1 lists all the attributes obtained from Quincy and Beutenberg weather networks.
@@ 4.2. Implementation of the Digital Twin Layer
We deploy the digital twin layer on NVIDIA Jetson Orin Developer Kit (12-core Arm Cortex-A78AE 64-bit CPU, 64 GB LPDDR5 RAM, and 2048-core GPU with 64 Tensor Cores) (karumbunathan2022nvidia, ). The algorithms are implemented in Python and executed on Linux (kernel version 5.15.148). We crawl Quincy weather station data from ZENTRA Cloud through HTTP requests using the Python requests package. The Beutenberg dataset is available as a .csv format, which Cerealia parses and stores locally in its internal database. We use TensorFlow (abadi2016tensorflow, ) for model training and inference purposes.

### v2

@@ 1. Introduction
Modern agriculture, especially high-value specialty crop production, relies heavily on data-driven automation technologies for production management. Agricultural decision-making increasingly relies on precise and consistent environmental data to optimize crop health, resource utilization, and yield predictions. Regional as well as localized weather critically drives several on-farm decision-support models and intelligent automated systems, including irrigation, abiotic stress (e.g., cold, heat) mitigation, and pest management, among others (Boomgard-Zagrodnik and Brown, 2022). Weather data measurements, which include essential metrics such as temperature, humidity, and precipitation, form the foundation of such decisions. However, weather networks, like any cyber-physical and Internet-of-Things (IoT) systems, are often plagued by inconsistencies [^1] In this work, we focus on effects on inconsistent/irregular weather streams rather than the causes that trigger the anomalies. due to sensor faults, calibration drift, communication errors, environmental interference, or cyber breaches. Imperfect measurements can significantly undermine the reliability of downstream applications in agriculture, ranging from predictive modeling to real-time interventions. Such discrepancies can lead to inaccurate predictions, suboptimal resource allocation, and reduced agricultural productivity.
Any inconsistent weather data input could lead to inadequate decision-making and wrongful actuation of precision management scenarios, resulting in considerable crop yield and quality losses. To address this problem, we adopt the digital twin technology (Barricelli et al., 2019) for analyzing and understanding agricultural weather networks that are critical for grower decision support. We build dynamic virtual data streams that mirror physical stations—to monitor, validate, and enhance the reliability of weather data in the face of anomalies. While digital twin technology is used in other cyber-physical domains (e.g., control systems, manufacturing, robotics), its benefits are not fully explored for digital agriculture. By leveraging the concept of digital twins, we show that it is feasible to create a virtual replica of weather networks for real-time inconsistency detection, data imputation, and predictive analytics. We name our framework Cerealia. [^2] In ancient Roman religion, the Cerealia was a major festival dedicated to the sowing, growing, and harvesting of crops. As our work aims to assist in better decision-making for agricultural stakeholders, we name our framework Cerealia.
@@
Cerealia is implemented and tested by using a commercial orchard located in Central Washington, United States (§4–§5). [^3] This research uses weather stations installed in a commercial orchard maintained by Washington State University AgWeatherNet team (46). While this is a closed system, access can be made available for research purposes by request, as we did in this work. We use this platform to demonstrate the efficacy of our approach in a real, live setup. Our use of this platform, however, by no means implies our affiliation with this infrastructure/institution. The digital twin modules are deployed on an NVIDIA Jetson Orin platform (Karumbunathan, 2022). In addition to live weather feed from the orchard, Cerealia is also tested using a publicly available European weather dataset (25). We conduct a comprehensive evaluation of nine state-of-the-art neural network models to analyze their effectiveness in detecting weather data inconsistencies. We measure the overheads of running learning models on our evaluation platform and test the scalability of Cerealia. We find that the overheads of checking (in terms of timing and memory usage) are minimal (less than 1 s and 0.5 MB of memory use). Our research highlights the potential of digital twin tools like Cerealia to improve agricultural decision-making under imperfect conditions.
Cerealia implementation is publicly available: https://github.com/CPS2RL/ag-dt.
@@ 2.1. Weather Data Impacts on Agriculture
Weather data has a significant impact on agriculture. For example, in early spring and late winter, frost damage in fruit crops can hurt fruit production. Fruit growers attempt to mitigate frost using wind machines, propane heaters, and under-tree sprinklers (Evans, 1999). To activate those mitigation tools, growers use weather data-driven models and sensing systems that automate and improve frost management with resource conservation (natural gas/energy) (Talsma et al., 2023; Cann et al., 2022). Similar to frost, extreme and more frequent heat stress events in summer can cause fruit cullage due to heat stress instigated by sunburn and related damages. To address this problem, researchers have explored weather-driven models to increase the effectiveness of heat stress mitigation techniques (Goosman et al., 2023; Ranjan et al., 2020; Amogi et al., 2023). In addition to frost and heat management, uninterrupted and reliable weather data obtained at the farm scale is necessary for several real-time tasks, including irrigation, disease and pest management, fruit growth prediction, among others (Boomgard-Zagrodnik and Brown, 2022).
Existing weather-driven agriculture decision-making models (Talsma et al., 2023; Cann et al., 2022; Goosman et al., 2023; Ranjan et al., 2020; Saxena et al., 2023; Amogi et al., 2023, 2024) rely on the availability of “perfect” weather information. Like any networked cyber-physical system, weather stations are also susceptible to faults and cyberattacks (Boomgard-Zagrodnik and Brown, 2022; Weiwei et al., 2018; Botschner et al., 2024; Kulkarni et al., 2024). Inadequate and inconsistent environmental data, for instance, due to faults in weather station sensors or adversarial breaches, often fail to protect the crop during extreme heat or cold events. Further, any unreliable decision-making model or lack of automation requires the growers to work on the farm in extreme weather conditions to manually activate protection measures (e.g., turning on heathers, putting protective netting, or tuning on the cooling system to prevent heat stress). We envision that a digital twin platform incorporating measurement inconsistencies provides a virtual view of the weather network and can help growers make informed decisions in their target agriculture applications.
@@ 2.2. Digital Twin and its Usefulness in Agriculture
In literature, a digital twin is defined as “a digital replica of a living or non-living physical entity” (El Saddik, 2018). Digital twin technology is used to describe and model the corresponding physical counterpart digitally by combining modeling, simulation, and emulation technologies with other analytics to better understand aspects of the target application domain. Digital twins are commonly used in many cyber-physical domains such as manufacturing (Kritzinger et al., 2018), IoT (Minerva et al., 2020), healthcare (Sun et al., 2023), and recently, digital agriculture (Purcell and Neubauer, 2023). However, its efficacy in weather-driven agricultural decision-making still has not been explored.
@@ 3.1.1. Physical Layer
As shown in Block 1 of Fig. 2, the physical layer consists of weather stations deployed in the field. Many off-the-shelf weather stations (e.g., Cabled Vantage Pro2 (10), ATMOS 41W (5)) and existing weather networks (e.g., AgWeatherNet (46)) generally collect field measurements from stations in a periodic interval and store them in a cloud gateway for further processing. We follow a similar approach in Cerealia, where a cloud storage service collects data from distributed stations and feeds this to our analysis engines in the digital twin layer (see next).
@@ 3.1.2. Digital Twin Layer
The key components of the digital twin layer (i.e., Blocks  3- 7) are modular by design. As mentioned above, our current setup incorporates four kinds of noisy data and uses nine machine-learning models to see the models’ behavior under these inconsistencies to find the best possible solution for Cerealia. However, other faulty/anomalous data and other statistical models or anomaly detection tools can be incorporated with Cerealia to check for inconsistencies. Likewise, as discussed in §6, for demonstration purposes, we tested Cerealia for two agricultural use cases (e.g., weather data imputation and fruit surface temperature prediction problems). However, this component of Cerealia (i.e., Block  7) can be adapted for any other agriculture decision-making tasks (e.g., fruit bloom phenology (Chaves et al., 2015), soil water content prediction (Nath et al., 2024)) without loss of generality.
@@ 3.2. Generating Noisy Measurements
One of the benefits of having a digital twin architecture like Cerealia is that we can emulate the behaviors of actual weather stations with customized parameters. Although physical station measurements could be anomalous in practice, (a) they are unpredictable, and (b) we need a reference model (i.e., ground truth) to check for abnormalities. Hence, we use a noise generator (De Bruijn et al., 2016) to mimic inconsistent behaviors. We consider four kinds of inconsistent data: (a) random (random spikes in the measurements), (b) malfunction (abrupt oscillations in the measurements), (c) drift (similar to malfunction, but the results are significantly shifted from nominal values), and (d) bias (results are scaled up or down than normal values). Figure 3 illustrates these four kinds of faults. Although we choose these four kinds of inconsistencies for our current demonstration—as they often represent typical sensor faults (Ni et al., 2009)—any other fault model can be integrated with Cerealia’s synthetic data generator module (i.e., Block 3 in Fig. 2), which then will be fed into the inconsistency checker (i.e., Block 5 in Fig. 2, discussed next).
@@ 3.3. Leveraging Machine Learning for Inconsistency Checking
Based on historical traces and a given noise model, we train machine learning models to characterize off-nominal data from real-time weather measurements generated from the physical layer. Some machine learning models, such as SVM (Suthaharan and Suthaharan, 2016), KNN (Steinbach and Tan, 2009), and K-Means (Steinley, 2006), often assume static data relationships or require handcrafted features, which may not be efficient for capturing the complex sequential nature of weather measurements. Also, they often struggle with large datasets, particularly high-dimensional and noisy data (which is the case for weather networks). Hence, we use deep learning models in Cerealia to check for inconsistencies. In our current implementation, we tested nine neural network models, which are discussed next. However, as mentioned earlier, other statistical or learning models can be integrated with Cerealia.
@@ 3.3.1. Convolutional Neural Network (CNN) Models
CNN models stand out in retrieving hierarchical and local features from time-series data like weather measurements. We use two CNN models: (a) Temporal Convolution Network (TCN) (He and Zhao, 2019) and (b) ResNet (He et al., 2020). TCN uses causal convolution and dilation to capture short-term and long-term temporal dependencies. ResNet uses residual connections and expedites the training of deep networks with reduced training error, but often requires careful tuning due to its complexity.
@@ 3.3.2. Recurrent Neural Network (RNN) Models
RNNs are highly effective for sequential data analysis due to their inherent memory capabilities. RNN models can capture the features and dependencies across long sequences. We incorporate the following RNN models for inconsistency checking: (a) Long Short-Term Memory (LSTM) (Van Houdt et al., 2020), (b) Bi-LSTM (Siami-Namini et al., 2019) and (c) Gated Recurrent Unit (GRU) (Shewalkar et al., 2019). LSTM is a well-known algorithm for retaining information over time. The latter two models are improved variants of LSTM that enhance sequence prediction.
@@ 3.3.3. Transformer Models
Transformer architectures are useful in tracking relationships in sequential data. We use two transformer-based models: (a) Time-Series Transformer (TST) (Zerveas et al., 2021) and (b) Informer (Zhou et al., 2021) that are useful for capturing intricate patterns and long sequences in the weather station measurements.
@@ 3.3.4. Hybrid Models
We also use two hybrid models: (a) TST-AE (Zhang et al., 2021) and (b) LSTM-AE (Provotar et al., 2019). TST-AE uses a transformer-based self-attention with an encoder-decoder structure for better feature extraction. LSTM-AE is based on an LSTM-based encoder-decoder and can provide efficient memory usage, but comes with the cost of reduced accuracy.
@@ 4.1. Physical Layer Integration and Real-Time Data Collection
We integrate Cerealia with a commercial apple orchard located in Quincy, Washington, United States ($47^{\circ}13^{\prime}31.3^{\prime\prime}$ $119^{\circ}57^{\prime}38.0^{\prime\prime}\text{W}$). The orchard has an average elevation of 322 m and an average slope of $2^{\circ}$ facing east. The orchard’s weather stations are maintained by the Washington State University AgWeatherNet (46) team. [^5] Note: The Majority of commercial agricultural weather networks are proprietary. Access to orchards/weather data is non-trivial, and few are publicly accessible for research. The station where we have access is installed at 1.5 m above ground level. The station uses an all-in-one weather sensor (ATMOS 41 from METER Group (5)). It measures several weather parameters such as solar radiation, precipitation, vapor pressure, relative humidity, barometric pressure, horizontal wind speed, wind gust, wind direction, tilt, lightning strike, and lightning average distance.
Weather stations sample measurements at 5-minute intervals and send the data to a vendor cloud service (ZENTRA Cloud (49)). We obtained weather data from March 2023 to January 2025. The collection frequency is twelve data tuples per hour per station (recall: stations send data to the cloud in 5-minute intervals), resulting in 192,613 data points. Following the naming convention suggested by the AgWeatherNet team, we refer to the measurements obtained from the stations as the “Quincy” network data. Our physical layer implementation demonstrates the feasibility of integrating Cerealia with a live operational weather network.
@@ 4.1.1. Preexisting Dataset
In addition to the live feed obtained from the orchard (Quincy network), we also use a publicly available dataset (named “Beutenberg”) (25). This weather data was recorded at the weather stations of the Max Planck Institute for Biogeochemistry in Jena, Germany. The dataset contains 20 different features, such as air temperature, atmospheric pressure, and humidity, between December 2015 to December 2020. The sampling rate for this dataset is six data points per hour. The total number of data points is 266,609. The reason for selecting this dataset is to test our inconsistency classifiers for diverse weather conditions (North America and Europe). Table 7 in Appendix A.1 lists all the attributes obtained from Quincy and Beutenberg weather networks.
@@ 4.2. Implementation of the Digital Twin Layer
We deploy the digital twin layer on NVIDIA Jetson Orin Developer Kit (12-core Arm Cortex-A78AE 64-bit CPU, 64 GB LPDDR5 RAM, and 2048-core GPU with 64 Tensor Cores) (Karumbunathan, 2022). The algorithms are implemented in Python and executed on Linux (kernel version 5.15.148). We crawl Quincy weather station data from ZENTRA Cloud through HTTP requests using the Python requests package. The Beutenberg dataset is available as a .csv format, which Cerealia parses and stores locally in its internal database. We use TensorFlow (Abadi et al., 2016) for model training and inference purposes.